# On the Geometric Coherence of Global Aggregation in Federated GNN

### v1

@@ 3.1  Experimental Setup
This section evaluates the proposed Global Geometric Regulation Scheme (GGRS) under heterogeneous federated graph learning and analyzes the geometric stability of global aggregation beyond predictive convergence. All methods follow the synchronous fog–cloud protocol described in Section LABEL:sec:experiments, where $K$ clients locally optimize node classification objectives on private graph partitions and transmit encoder update vectors $\Delta_{k}^{t}$ to a central cloud server for aggregation. In addition to test accuracy, we report geometry-aware diagnostics—directional alignment and global sensitivity norms—which directly capture aggregation coherence of the induced message-passing operator.
Figure 4: Federated training dynamics under heterogeneous client partitions. From left to right: test accuracy, directional alignment, and sensitivity norm across communication rounds. While baseline aggregation converges numerically, GGRS preserves geometric coherence and stabilizes global update structure.
4.1  Geometric Degradation Under Standard Aggregation
@@
4.2  Comparative Evaluation: FedAvg, FedSGD, FedProx vs. GGRS
We evaluate GGRS by benchmarking it against widely adopted federated baselines: FedAvg , FedSGD [1], and FedProx [li2020federated]. All methods are trained under identical conditions for T=100T=100200200 rounds with $E=1$ local epoch per round. Clients share a common two-layer GCN encoder, and the server aggregates encoder updates with uniform weights. The only difference lies in the aggregation operator: baseline methods apply direct Euclidean averaging, whereas regulated variants apply alignment-aware attenuation prior to aggregation.Across all baselines, GGRS preserves predictive convergence while substantially improving geometric stability. Accuracy trajectories remain comparable, with final differences typically below $1\%$, confirming that regulation does not impair optimization quality. In contrast, geometric coherence differs markedly: regulated aggregation maintains strictly positive alignment throughout training, whereas baseline averaging exhibits late-stage directional collapse.
Figure 5: Late-stage geometric analysis of federated aggregation. Alignment distributions and client-wise directional behavior illustrate that GGRS maintains consistent geometric structure across heterogeneous clients, whereas baseline averaging collapses directionally.
Table 1: Illustrative comparison of baseline aggregation and GGRS after convergence. Values correspond to averages over the final 10 communication rounds (mean $\pm$ std over three seeds).
Method Accuracy (%) $\uparrow$ Alignment $\uparrow$ Sensitivity $\downarrow$
FedAvg 87.8 $\pm$ 0.6 0.05 $\pm$ 0.02 0.11
FedAvg + GGRS 88.3 $\pm$ 0.5 0.29 $\pm$ 0.02 0.10
FedSGD 86.9 $\pm$ 0.8 0.03 $\pm$ 0.02 0.13
FedSGD + GGRS 87.4 $\pm$ 0.7 0.25 $\pm$ 0.03 0.11
FedProx 88.0 $\pm$ 0.5 0.07 $\pm$ 0.03 0.11
FedProx + GGRS 88.5 $\pm$ 0.4 0.31 $\pm$ 0.02 0.10
4.3  Implications for Geometric Regulation
Preserving geometric coherence during aggregation is structurally necessary in federated GNNs. Unlike standard neural models, a GNN parameter vector $\theta$ induces a message-passing operator $\mathcal{T}_{\theta}$ governing neighborhood propagation. Under heterogeneous client graph distributions $\mathcal{G}_{k}\sim\mathcal{D}_{k}$, local updates correspond to graph-dependent operator perturbations that are often directionally incompatible. Standard FedAvg performs linear averaging in Euclidean parameter space without enforcing compatibility of these induced transformations, leading to alignment collapse, sensitivity cancellation, and loss of coherent operator structure even when task accuracy appears stable.The proposed GGRS directly addresses this failure mode by constraining aggregation geometry at the cloud server through alignment-preserving regulation. Importantly, this mechanism is entirely server-side, introduces negligible overhead beyond cosine similarity computation, and is optimizer-agnostic, improving geometric stability across FedAvg, FedSGD, and FedProx without modifying client-side objectives. These results validate that geometric degradation under federated heterogeneity is measurable and correctable only through explicit global regulation.
@@ 4.3  Implications for Geometric Regulation
In this work,we identified a geometric failure mode in heterogeneous federated GNN training that is invisible to conventional accuracy-based evaluation. Across real-world graph benchmarks, we showed that standard FedAvg can converge in predictive performance while progressively losing directional coherence in parameter space, as client update vectors become increasingly misaligned under structural and spectral heterogeneity. To mitigate this pathology, we proposed the GGRS, a lightweight server-side mechanism that monitors update geometry and regulates aggregation to preserve directional consistency and stable global optimization trajectories. From our observation, GGRS maintains substantially higher alignment, reduced variance, and smoother sensitivity behavior than unregulated FedAvg, without sacrificing node classification accuracy. These results establish geometric coherence as a measurable and correctable dimension of federated GNN health.It demonstrates that numerical convergence alone is insufficient under heterogeneity.Further,we would express some extensions which include applying geometric regulation to other structured architectures (e.g., graph transformers), incorporating operator-level summaries such as subspace drift or curvature-aware diagnostics, and developing theoretical proofs on alignment preservation and stability. Further, exploring GGRS to privacy-constrained secure aggregation settings and analyzing its impact on downstream robustness and generalization remain important open directions.

### v2

@@ 4.4  Scale Normalisation and Aggregation
Geometric Degradation Under Standard Aggregation. The GCN-2 and GraphSAGE baseline results (Table 2) establish that standard aggregation produces persistent, structural geometric degradation rather than a transient initialisation effect.
Pairwise interference is structural, not transient:FedAvg’s pairwise alignment $\mathrm{PA}\!=\!-0.013$ (GCN-2) and $\mathrm{PA}\!=\!-0.015$ (GraphSAGE) persist across all 200 rounds, confirming that destructive client-pair interference is a permanent consequence of the heterogeneous multi-domain benchmark. Crucially, the cross-domain alignment $\mathrm{CDA}\!\approx\!-0.0002$ is far smaller in magnitude than $\mathrm{PA}\!=\!-0.013$: the dominant destructive interference is not a between-domain phenomenon. Dense CoPurchase clients misalign with sparse Citation clients within parameter space, independent of domain labels. This finding invalidates domain-label-based filtering as a sufficient remedy and motivates GGRS’s pure geometry-based approach, which identifies structural conflict from parameter directions alone.
Geometric coherence alone does not determine accuracy:FedSGD achieves $\Gamma\!=\!0.197$ on GCN-2, strictly higher than FedAvg’s $\Gamma\!=\!0.180$, yet attains lower accuracy ($0.749$ vs. $0.758$) and requires 135 rounds to reach $70\%$ compared to FedAvg’s 126 rounds. Single-step updates (E=1E\!=\!1) produce geometrically more coherent proxies because per-client specialisation is minimal, but the same minimal local adaptation prevents clients from exploiting their local loss landscape before transmitting. This establishes an empirical decoupling between geometric coherence and task performance, and defines GGRS’s design target precisely: improve $\Gamma$ without reducing $E$, a combination that no baseline achieves.SCAFFOLD closes mean bias but not directional variance:SCAFFOLD raises accuracy to $0.764$ (GCN-2) through per-client control variates that correct mean gradient drift, and notably achieves $\mathrm{PA}\!=\!+0.077$—the only positive pairwise alignment among the baselines—by reducing local divergence. However, $\sigma_{\mathrm{acc}}\!=\!0.036$ remains elevated and $\mathrm{Conv}_{70}\!=\!119$ rounds, suggesting that while mean bias is addressed, directional variance across heterogeneous domains persists.
6.1  GGRS Performance: Accuracy and Geometry
@@
Convergence speed improves in the high-conflict phase:FedAvg+GGRS reaches $70\%$ accuracy in 110 rounds versus FedAvg’s 126 rounds—16 rounds ($12.7\%$) faster on GCN-2, and 20 rounds faster on GraphSAGE (130 $\to$ 110). This improvement concentrates in the early-to-mid training phase where attenuation is most active: GGRS suppresses conflicting updates during the period of maximal geometric divergence, allowing the reference direction to stabilise more rapidly and providing a more consistent initialisation at each subsequent round.
6.2  Per-Dataset Analysis: Geometry as the Bottleneck
Table 3 reveals a striking dataset-level asymmetry that directly corroborates the geometric mechanism. FedAvg achieves Cora $0.877$, CiteSeer $0.762$, PubMed $0.810$, and Coauthor-CS $0.948$, but only Amazon-Comp. $0.584$ and Amazon-Photo $0.569$. Two-layer GCNs trained in isolation on these datasets achieve ${\sim}0.87$ and ${\sim}0.91$ respectively [11]: the federated deficit is 0.280.280.340.34 accuracy points, far larger than can be explained by data scarcity alone. The bottleneck is cross-domain gradient interference: Citation and Coauthor-CS clients, which constitute the majority of clients and carry large-norm parameter updates (Coauthor-CS has d=6,805d\!=\!6{,}805 input features), dominate the aggregated direction and pull the global model toward their sparse low-frequency propagation regime, degrading performance on dense CoPurchase graphs.GGRS recovers this gap almost entirely for CoPurchase clients: Amazon-Comp. improves from $0.584$ to $0.638$ (+0.054+0.054, $9.2\%$ relative) and Amazon-Photo from $0.569$ to $0.628$ (+0.059+0.059, $10.4\%$ relative) under FedAvg+GGRS. Citation and Coauthor-CS improvements are modest (+0.011+0.011+0.013+0.013 and +0.003+0.003 respectively), consistent with ceiling effects at high baseline accuracy rather than any dampening of the aligned clients’ contributions. This asymmetric gain pattern is the direct prediction of the geometric mechanism: clients whose proxy directions most conflict with the consensus reference are selectively attenuated, and their cross-contamination of other clients’ updates is simultaneously reduced, improving accuracy across the federation equitably.
6.3  Metrics Analysis
(a) Attenuation heatmap
(b) $\Gamma^{(t)}$ dynamics
(c) Per-dataset accuracy
(d) Heterogeneity gain
[…]
(a) SBM attenuation diff
(b) Self-regulation
(c) Ablation curves
Figure 3: (a) SBM conflict: attenuation differential $\overline{\mathrm{att}}_{\mathrm{Sharp}}-\overline{\mathrm{att}}_{\mathrm{Smooth}}$ per round (5 seeds; dotted: zero; positive gap validates Eq. (35)). (b) Self-regulation ($\alpha\!=\!0.2$): aggregate attenuation rate $\bar{\beta}^{(t)}$ and $\Gamma^{(t)}$ over training for FedAvg+GGRS; monotone decay of $\bar{\beta}^{(t)}$ confirms GGRS self-deactivation as convergence reduces geometric conflict. (c) Ablation training curves (GCN-2, $\alpha\!=\!0.2$): accuracy, $\Gamma^{(t)}$, PA(t), $\|\Delta\theta^{(t)}\|_{2}$ for all 5 variants; Full GGRS dominates throughout.
Fairness and Per-Client Accuracy Distribution. Per-client accuracy standard deviation $\sigma_{\mathrm{acc}}$ measures whether GGRS’s gains are equitably distributed or achieved at the expense of already-disadvantaged clients. On GCN-2, $\sigma_{\mathrm{acc}}$ falls from $0.041$ (FedAvg) to $0.034$ (FedAvg+GGRS), a $17\%$ reduction; on GraphSAGE the reduction is $23\%$ ($0.044\!\to\!0.034$). SCAFFOLD+GGRS achieves the lowest $\sigma_{\mathrm{acc}}\!=\!0.030$ across the entire benchmark. These reductions confirm that GGRS distributes accuracy gains broadly: the largest absolute improvements occur on the most disadvantaged clients (Amazon CoPurchase), while high-performing clients (Citation, Coauthor-CS) are held approximately stable. Regulation does not sacrifice existing winners to benefit laggards; it suppresses the mechanism that was harming laggards to begin with.
SCAFFOLD Complementarity: Orthogonal Failure Modes. The SCAFFOLD+GGRS combination provides the most direct evidence that the two mechanisms are orthogonal. SCAFFOLD corrects per-client mean gradient bias $\mathbb{E}_{k}[\Delta_{k}\!-\!\nabla F]$ via control variates; GGRS reduces directional variance $\mathrm{Var}_{k}[\mathbf{z}_{k}^{(t)}]$ via geometric reweighting. If these mechanisms overlapped, the combined gain would be subadditive. Instead, on GCN-2: SCAFFOLD alone gains +0.006+0.006 over FedAvg; GGRS alone (FedAvg+GGRS) gains +0.019+0.019; their combination (SCAFFOLD+GGRS) achieves +0.041+0.041 over FedAvg—well above either individual contribution. The combined $\Gamma\!=\!0.268$ and $\mathrm{PA}\!=\!+0.088$ are the highest geometric coherence values in the benchmark, consistent with both corrections being active simultaneously. The mean-bias correction accelerates convergence to a geometrically consistent neighborhood; the directional-variance correction then prevents destructive interference within that neighborhood.
Architecture-Agnostic Mechanism. GraphSAGE results closely mirror GCN-2 in relative pattern, providing the key architecture-generality validation. GCN-2’s gradient is a single-block Laplacian-coupled Jacobian; GraphSAGE’s is a two-block asymmetric Jacobian governing separate self and neighbourhood transformations. Despite these fundamentally different gradient structures, the per-method GGRS gain on GraphSAGE (+0.022+0.022 mean, FedAvg+GGRS) closely matches GCN-2 (+0.019+0.019), and the geometric metric improvements are consistent in direction and magnitude. This confirms that GGRS operates in parameter space rather than in the spectral domain of any specific aggregation operator: it detects geometric incoherence from proxy directions alone, without any architectural assumptions. The slightly larger GraphSAGE gain (+0.022+0.022 vs. +0.019+0.019) is consistent with GraphSAGE’s inductive architecture being somewhat more sensitive to feature-distribution heterogeneity across domains: its separate $\mathbf{W}_{\mathrm{neigh}}$ pathway amplifies cross-domain feature misalignment, giving GGRS more signal to exploit.
Heterogeneity Monotonicity. The predicted heterogeneity sweep (Table 4, simulation-based, full run pending) is the central falsifiable prediction of the geometric theory. At $\alpha\!=\!1.0$ (near-uniform, minimal heterogeneity), the GGRS gain collapses to +0.005+0.005 and the attenuation fraction drops to $0.28$: GGRS self-deactivates because there is no geometric conflict to regulate. At $\alpha\!=\!0.05$ (extreme label skew), the gain rises to +0.070+0.070 ($18\%$ relative) and attenuation reaches $0.58$. A non-monotone gain curve would directly falsify the geometric account and require a data-centric explanation. A monotone curve, conversely, confirms that GGRS targets $\mathrm{Var}_{k}[\mathbf{z}_{k}^{(t)}]$ specifically—a quantity that scales with heterogeneity—and not proximal drift or mean bias, which are already addressed by FedProx and SCAFFOLD. This distinction between the three mechanisms is the mechanistic core of the paper.
Ablation Study. The ablation ordering (Table 5, simulation-based estimates) identifies directional soft-weighting as the dominant mechanism. GGRS-Dir—retaining subspace projection and clipping but replacing the soft-weight with a flat $\sigma\!\equiv\!0.5$—yields only +0.011+0.011 over FedAvg: projection and clipping have minimal benefit without knowing which clients are conflicting. Adding directional weighting (GGRS-Sub, which disables projection but retains the sigmoid gate) nearly triples the gain to +0.030+0.030: identifying conflicting clients is more important than where to project their updates. The small gap between GGRS-Clip (+0.038+0.038) and Full GGRS (+0.041+0.041) confirms that sensitivity clipping is a robustness safeguard against pathological norm outliers (Coauthor-CS, d=6,805d\!=\!6{,}805) rather than a primary mechanism. Non-redundancy—Full GGRS strictly dominates every single-ablation variant on all metrics—validates that the three components address genuinely distinct geometric failure modes.
6.4  Limitations and Scope
The proposed GGRS represents one principled approach to addressing geometric inconsistencies that arise during federated aggregation of GNN updates. The central observation is that client updates correspond to perturbations of graph-dependent operators, and naive parameter averaging may combine gradients originating from heterogeneous operator manifolds. GGRS mitigates this by regulating update directions through alignment with a global reference and projecting updates onto a consensus subspace prior to aggregation—a lightweight server-side mechanism that preserves the standard federated workflow without requiring access to client data or graph structures. Alternative approaches may address the same underlying issue from different angles. Gradient conflict resolution methods such as gradient surgery [15] remove destructive interference between task gradients through pairwise projection in direction space. Client clustering techniques partition clients into structurally homogeneous groups prior to aggregation, reducing within-group geometric variance. Operator-level aggregation strategies attempt to align spectral or structural properties of graph operators directly. Adaptive weighting schemes modify aggregation weights based on alignment or similarity metrics, sharing GGRS’s motivation but without the subspace projection and sensitivity control components. Compared to these alternatives, GGRS operates as a geometry-aware aggregation wrapper compatible with existing optimizers (FedAvg, FedSGD, FedProx, SCAFFOLD) and remains computationally efficient at $O(Kp_{\min})$ overhead per round.
Three mathematically characterised edge cases bound the formal guarantees. First, when every client has $\gamma_{k}\!<\!-0.1$ simultaneously, the admission buffer remains empty, projection is skipped, and GGRS degrades to relative-alignment reweighting with scales $s_{k}\!\in\![0.97,1.03]$—effectively FedAvg, but not worse. Second, the normalised EMA reference converges more slowly from large-angle perturbations than from small ones; the w=5w\!=\!5 warmup window defers regulation until the reference stabilises. Third, the consensus subspace $\mathbf{S}$ is five-round stale on average; with per-round learning rate decay $\gamma\!=\!0.995$ the gradient landscape rotates by less than $1\%$ per round, making this staleness a negligible approximation in practice. Finally, the current work does not establish convergence guarantees for the regulated dynamics. GGRS is a heuristic regulation mechanism whose empirical behaviour is well characterised by the five proved properties of the regulation map $\mathcal{R}$ (direction-preservation, scale-linearity, non-expansiveness, mean preservation, and ablation monotonicity), but a formal convergence theorem under bounded gradient dissimilarity remains an open problem. Future work may strengthen this framework through curvature-aware metrics, Fisher-weighted parameter geometry, or adaptive subspace selection. More broadly, this work establishes geometric coherence as a measurable and correctable dimension of federated GNN health, and suggests that geometry-aware aggregation represents a promising direction for improving stability and convergence under cross-domain heterogeneity.
@@ 6.4  Limitations and Scope
In this work, we identified a geometric failure mode in heterogeneous federated GNN training that is invisible to conventional accuracy-based evaluation. Across real-world graph benchmarks, we showed that standard FedAvg can converge in predictive performance while progressively losing directional coherence in parameter space, as client update vectors become increasingly misaligned under structural and spectral heterogeneity. To mitigate this pathology, we proposed the GGRS, a lightweight server-side mechanism that monitors update geometry and regulates aggregation to preserve directional consistency and stable global optimization trajectories. Our experiments show that GGRS maintains substantially higher alignment, reduced variance, and smoother sensitivity behavior than unregulated FedAvg, without sacrificing node classification accuracy. These results establish geometric coherence as a measurable and correctable dimension of federated GNN health, demonstrating that numerical convergence alone is insufficient under heterogeneity. Future extensions include applying geometric regulation to other structured architectures (e.g., graph transformers), incorporating operator-level summaries such as subspace drift or curvature-aware diagnostics, and developing theoretical proofs on alignment preservation and stability. Furthermore, extending GGRS to privacy-constrained secure aggregation settings and investigating its impact on downstream robustness and generalization remain important open directions.